a first order parallel reaction given as"a-----b and c .given that k1 is rate constant for formation of b and k2 is rate constant for c.k1=1.3 * 10 to power -5 ,sec to power -1.also k1\k2=1\9.find conc.ratio of c to a if an experiment is allowed to start with only a for 1 hr. I HAVE ALREADY ASKED ANS.FOR THIS Q.THE ANS GIVEN IS 0.537.THE ANSWER U GAVE IS 9/11.PLZ GIVE ME A RIGHT SOLUTION.


 




 

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Aditya,

Hi,

A---->B     k1

A---->C     k2

d[B]/dt = k1*[A]      d[B] = k1*[A] dt

d[C]/dt = k2*[A]      d[C] = k2*[A] dt

therefore [B]/[C] = k1/k2 = 1/9

[A]0 = [A]+[B]+[C]

[A]=[A]0e-(k1+k2)t

[A]=[A]0e-0.468

[A]0 = [A]e0.468

[A]0 = [A]+[B]+[C]

[A]e0.468= [A]+[B]+[C]

[A]*1.656= [A]+[B]+[C]

[A]*0.656=[B]+[C]

0.656*[A]/[C] = [B]/[C] + 1 = 1/9 +1 = 11/9

[C]/[A]=0.656*9/11=0.537